### v2

@@ 1 INTRODUCTION
The paper is set up as follows: Section 1 serves as a comprehensive foundation, beginning with the definition of the fill management problem and the establishment of the study’s context, followed by a review of relevant literature and various approaches within the field. It then transitions into a discussion of our research methodologies, encompassing a range of techniques and methods employed in our investigation. In Section 2, we detail our proposed system’s design and architecture. This section also introduces our Sonic Impulse Response Ensemble Classifier (SIREC) algorithm and describes its concept and operational details. Section 3 is dedicated to presenting our findings, where we validate our results through various scenarios. The paper concludes in Section 4 with a discussion that synthesizes our main insights and reflects on the study’s implications.
@@ 1.1 RELATED WORK
1.2 SYSTEM OVERVIEW
This project focuses on the development of an IWMS that utilizes on-device AI for the classification of waste container fill levels. As a proof of concept, the inner side of a waste container’s lid was equipped with an ESP8266, a compact Wi-Fi-capable micro-controller [19], to which a passive buzzer and a microphone were attached, as shown in Figure 1. The ESP8266 board (4) served as the central processing unit of the system, generating sound signals (sine sweeps) through the passive buzzer (2) and capturing the resulting acoustic response (bin interior) via the microphone (3). The captured and preprocessed (see Section 2) signal was then transmitted to a Raspberry Pi for storage using MQTT [20], a lightweight messaging protocol designed for small sensors and mobile devices.
@@ 1.3 SPECTRAL SUBTRACTION
1.4 SPATIAL IMPULSE RESPONSE
@@ 2 SYSTEM MODEL
2.1 DATA GENERATION & PREPROCESSING
Starting with the gray area in the presented flowchart, an ESP8266 utilizes pulse width modulation (PWM)–a method used to encode the amplitude of a signal into a duty cycle, which controls the power delivered to a load–to generate a linear sinusoidal sweep. This sweep serves as an auditory input signal produced via a passive buzzer. The signal is subsequently captured by a microphone, representing the output signal $y[n]$. Spectral subtraction is then employed for noise reduction of the output signal. This involves converting the signal into its frequency spectrum via FFT, followed by subtracting the estimated noise. This noise estimate is derived from a designated section recorded by the microphone where the sinus sweep is inaudible, hence, only background noise is present. The noise is transformed into its frequency spectrum, and then the disruptive frequencies are subtracted from the output signal $y[n]$. Post spectral subtraction, the RIR is approximated. Given that an impulse response uniquely identifies an LTI system, the most straightforward choice for a measuring signal to determine the RIR would be a Dirac or Kronecker impulse. An idealized impulse, such as the Dirac delta function in continuous systems, theoretically spans the entire frequency range, signifying that all frequencies are stimulated simultaneously and with equal intensity, as evidenced by its Fourier transform yielding a constant value across all frequencies [27]. Similarly, the Kronecker delta function [23] in discrete systems exhibits an analogous property, with its Fourier transform resulting in a constant value for all frequency indices $k$, limited to a range up to the so-called Nyquist frequency[25]. However, generating such idealized impulses with a sound source is impractical, for instance, due to the infinitely short duration of the signals[28]. Hence, alternative methods that mimic the properties of the ideal impulses but are feasible in reality are often used in electro-acoustic systems to stimulate sound impulses. The linear sinusoidal sweep is one such implementable signal[26]. It uses an excitation signal that covers as broad a frequency range as possible to enable a comprehensive measurement of a room’s acoustic properties. Equation 1 describes such a linear sinusoidal sweep that varies the frequency from $f_{0}$ to $f_{1}$ over time $T$, where $f_{s}$ denotes the sampling rate [29].
@@ 2.1 DATA GENERATION & PREPROCESSING
Figure 3: Employed sine sweep signal.
Algorithms 2 and 1 offer the pseudo-code for spectral subtraction and for the estimation of the RIR, respectively.
@@
2.2 TRAINING & TESTING
@@
Following the extraction of RIRs, the data is transmitted to a Raspberry Pi via MQTT. There, Python scripts manage and utilize the data for two primary processes, depicted as numbers 2 and 3 in Figure 2, which involve the SIREC algorithm developed for this project. During the training process, RIRs corresponding to varying fill levels of the waste container are collected and labeled. An ensemble ML model, specifically the SIREC model, is then trained using these RIRs, employing a predetermined number of estimators ($N_{\text{estim.}}$), which are DTs. For each tree, two intervals are generated: a random interval $I_{\text{rnd.}}$, whose length varies within a adjustable maximum and minimum range across the total length of the RIRs, and a fixed interval $I_{\text{fix.}}$, matching the length of the random interval but positioned at the start of the RIRs. Features are then extracted from both intervals to train the respective tree. Within $I_{\text{rnd.}}$, the magnitude of the frequency spectrum is determined using FFT, and the ratio of the minimum to maximum values is extracted. Empirical analysis has shown these features to be significant for classification. Within $I_{\text{fix.}}$, two feature classes are also extracted, but based on novel approaches, incorporating the mean and standard deviation as characteristic features of the time series into the tree’s training. Time series data can have varying structures, and in some cases, they may exhibit symmetry along the x-axis without a vertical shift. In this project, the RIR is calculated in a manner that closely exhibits these two properties. An example of an extracted RIR is shown in Figure 4. Considering this structure of the RIRs, calculating the mean would yield a value close to zero for different classes, making it also problematic for the standard deviation, which depends on the mean. In light of this, we propose an alternative calculation method for the fixed interval, computing the difference between each pair of adjacent values across all $N-1$ values to determine the mean and standard deviation of these newly generated $N-1$ values. This approach captures not only the magnitude of the mean and standard deviation but also the characteristic property of each RIR, combined with these two measures. The necessary computational steps are outlined in Equations 2 and 3.
@@
$\forall I_{\text{fix.}}\in\mathbb{R}^{n}\text{ with }I_{\text{fix.}}=[\iota_{1},\iota_{2},...,\iota_{\text{n}}]^{T}:$
$\mu_{\text{diff.}}=\frac{\sum_{k=1}^{n-1}(\iota_{\text{k}}-\iota_{\text{k+1}})}{n-1}$ (2)
$\sigma_{\text{diff.}}=\sqrt{\frac{1}{n-1}\sum_{k=1}^{n-1}\left(\iota_{\text{k}}-\iota_{\text{k+1}}-\mu_{\text{diff.}}\right)^{2}}$ (3)
Each of the $N_{\text{estim.}}$ trees is trained using this feature extraction methodology, forming an independent ML model that, in combination, constitutes SIREC, the complete ensemble ML model. In the testing process, depicted as number 3 in Figure 2, the model’s accuracy is validated using a portion of the collected RIRs and their corresponding fill level labels. Here, SIREC utilizes the intervals created during its training process, extracts relevant features, and aggregates predictions from all its DTs to derive a final result. The F1 score metric is then used to determine if the validation is sufficiently accurate to, if successful, convert the trained and validated SIREC model into C++ code. This code can then be seamlessly uploaded to the ESP8266 board using the Arduino IDE [31]–a development environment that facilitates programming and interfacing with various micro-controllers and modules. Thus, enabling local predictions using the trained ML model without the need to delegate the classification task to external resources.
3 VALIDATION
@@
3.1 PERFORMANCE
Data was systematically collected under varied conditions to benchmark different classification models. A selection of the characteristics of these data collection environments is presented in Table 1. The first column in the dataset uniquely identifies each entry. The next five columns present the distribution of data across designated categories. The final column indicates the type of filler material used, which is either straw or cardboard (CB). A linear sine-sweep signal (Figure 3), used to derive these measurements, was played between samples 2048 and 3072 of a 4096-sample-long recording captured at a sampling rate of 10kHz. The sample interval prior to this range was utilized for noise estimation, instrumental in noise attenuation via spectral subtraction.
Table 1: Extract of Accumulated Measurement Data
@@ 3.1 PERFORMANCE
It is particularly noteworthy that SIREC and the Random Interval Spectral Ensemble (RISE) [34] algorithm, which secured second place in Figure 5, share two fundamental commonalities. Both methodologies are predicated on an ensemble of DTs and both engage in the generation of random intervals, within which they extract the magnitude of the frequency spectrum as one of the features. However, there are key advantages that distinguish SIREC from RISE, rendering it more beneficial. Contrary to RISE, SIREC exhibits superior speed and energy efficiency (as will be discussed in the following section and shown in Table 3), as well as a significant improvement in memory efficiency. This is largely due to the fact that RISE computes the auto-correlation function, with the default setting being 100 auto-correlation terms, which inherently makes SIREC more suited for local operation on micro-controllers for classification tasks. Furthermore, SIREC tends to demonstrate marginally better average accuracy. To support this claim, Figure 6, as an extension to Figure 5, provides a detailed depiction of the F1 scores between RISE and SIREC over the 100 iterations.
@@
A significant advantage of the Random Interval Generation concept for feature extraction, as implemented in SIREC, is its inherent reduction or elimination of the need for expert knowledge about the collected data to train the model. For instance, a stochastic search across various randomly generated intervals within specified minimum and maximum lengths allows for the bypassing of such expert knowledge. This stochastic search method ensures the identification of the ideal data segment for feature extraction without necessitating extensive manual data analysis. Moreover, since SIREC is based on DTs, which represent a symbolic form of representation, acquiring expert knowledge about the data post-training is straightforward. Our stochastic search for optimal SIREC parameters (thus, the best random intervals) has achieved accuracies up to 100% in some cases. The results of this stochastic search, along with the parameters used for SIREC, are presented in Table 2.
Table 2: Stochastic Search for SIREC with Training Data $T_{1}$ $T_{2}$, Test Data $T_{3}$ $T_{4}$.
Segment Length N-Estim. Max. Interval Length Min. Interval Length Random State F1 Score
@@
3.2 ENERGY MEASUREMENTS
In the analysis of energy consumption, our investigation reveals significant insights into the trade-offs between computational energy efficiency and predictive accuracy across three distinct algorithms: our proposed algorithm (SIREC), Time Series Forest Classifier (TSFC)[35] and RISE. The measurement method and experiment setup is based upon our previous work [36, 37] and explained in detail by [38]. The resource measurements adhere to the methodology and guidelines outlined in the Green Software Measurement Model [39]. Measurements were carried out using a computer system that included an Intel Core i5-650 CPU. The system’s memory was composed of two 2 GB RAM modules, totaling 4 GB. For storage, the computer utilized a dual-setup, comprising a 500 GB hard disk drive (HDD) and a 250 GB solid-state drive (SSD). The analysis is structured around various scenarios–configurations of segment lengths and numbers of estimators, focusing on key metrics such as energy consumption during training and testing phases, computational time, CPU usage, and F1 score. These scenarios, along with their corresponding results listed in Table 3, were tested on the same datasets, $T_{1}$ and $T_{2}$ for training, and $T_{3}$ and $T_{4}$ for testing, as detailed in Table 1. It is important to highlight that a random state parameter was employed across all algorithms to ensure comparability and control over experimental variability. A value of 7 was selected arbitrarily for the random state, with no intent to favor any specific algorithm, especially not SIREC.

### v3

@@ I Introduction
The paper is set up as follows: Section I serves as a comprehensive foundation, beginning with the definition of the fill management problem and the establishment of the study’s context, followed by a review of relevant literature and various approaches within the field. It then transitions into a discussion of our research methodologies, encompassing a range of techniques and methods employed in our investigation. In Section II, we detail our proposed system’s design and architecture. This section also introduces our Sonic Impulse Response Ensemble Classifier (SIREC) algorithm and describes its concept and operational details. Section III is dedicated to presenting our findings, where we validate our results through various scenarios. The paper concludes in Section IV with a discussion that synthesizes our main insights and reflects on the study’s implications.
@@ I-A Related Work
I-B System Overview
This project focuses on the development of an IWMS that utilizes on-device AI for the classification of waste container fill levels. As a proof of concept, the inner side of a waste container’s lid was equipped with an ESP8266, a compact Wi-Fi-capable micro-controller [19], to which a passive buzzer and a microphone were attached, as shown in Figure 1. The ESP8266 board (4) served as the central processing unit of the system, generating sound signals (sine sweeps) through the passive buzzer (2) and capturing the resulting acoustic response (bin interior) via the microphone (3). The captured and preprocessed (see Section II) signal was then transmitted to a Raspberry Pi for storage using MQTT [20], a lightweight messaging protocol designed for small sensors and mobile devices.
@@ I-C Spectral Subtraction
I-D Spatial Impulse Response
@@ II System Model
II-A Data Generation & Preprocessing
Starting with the gray area in the presented flowchart, an ESP8266 utilizes pulse width modulation (PWM)–a method used to encode the amplitude of a signal into a duty cycle, which controls the power delivered to a load–to generate a linear sinusoidal sweep. This sweep serves as an auditory input signal produced via a passive buzzer. The signal is subsequently captured by a microphone, representing the output signal $y[n]$. Spectral subtraction is then employed for noise reduction of the output signal (see pseudo-code Algorithm 1). This involves converting the signal into its frequency spectrum via FFT, followed by subtracting the estimated noise. This noise estimate is derived from a designated section recorded by the microphone where the sinus sweep is inaudible, hence, only background noise is present. The noise is transformed into its frequency spectrum, and then the disruptive frequencies are subtracted from the output signal $y[n]$. Post spectral subtraction, the RIR is approximated (see pseudo-code Algorithm 2). Given that an impulse response uniquely identifies an LTI system, the most straightforward choice for a measuring signal to determine the RIR would be a Dirac or Kronecker impulse. An idealized impulse, such as the Dirac delta function in continuous systems, theoretically spans the entire frequency range, signifying that all frequencies are stimulated simultaneously and with equal intensity, as evidenced by its Fourier transform yielding a constant value across all frequencies [27]. Similarly, the Kronecker delta function [23] in discrete systems exhibits an analogous property, with its Fourier transform resulting in a constant value for all frequency indices $k$, limited to a range up to the so-called Nyquist frequency[25]. However, generating such idealized impulses with a sound source is impractical, for instance, due to the infinitely short duration of the signals[28]. Hence, alternative methods that mimic the properties of the ideal impulses but are feasible in reality are often used in electro-acoustic systems to stimulate sound impulses. The linear sinusoidal sweep is one such implementable signal[26]. It uses an excitation signal that covers as broad a frequency range as possible to enable a comprehensive measurement of a room’s acoustic properties. Equation 1 describes such a linear sinusoidal sweep that varies the frequency from $f_{0}$ to $f_{1}$ over time $T$, where $f_{s}$ denotes the sampling rate [29].
@@ II-A Data Generation & Preprocessing
Figure 3: Employed sine sweep signal.
II-B Training & Testing
Following the extraction of RIRs, the data is transmitted to a Raspberry Pi via MQTT. There, Python scripts manage and utilize the data for two primary processes, depicted as numbers 2 and 3 in Figure 2, which involve the SIREC algorithm developed for this project. During the training process, RIRs corresponding to varying fill levels of the waste container are collected and labeled. The collection of the RIRs can be represented by Equation (2):
$\mathbf{X}_{S,1:N}=\begin{bmatrix}v_{1,1}&v_{1,2}&\cdots&v_{1,p_{i}}&\cdots&v_{1,q_{i}}&\cdots&v_{1,N}\\ v_{2,1}&v_{2,2}&\cdots&v_{2,p_{i}}&\cdots&v_{2,q_{i}}&\cdots&v_{2,N}\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ v_{S,1}&v_{S,2}&\cdots&v_{S,p_{i}}&\cdots&v_{S,q_{i}}&\cdots&v_{S,N}\end{bmatrix}$ (2)
The corresponding labels are given by Equation (3):
$\mathbf{y}=\begin{bmatrix}l_{1},&l_{2},&\ldots,&l_{S}\end{bmatrix}^{T}$ (3)
where $\mathbf{X}_{S,1:N}\in\mathbb{R}^{S\times N}$, with $S$ as the number of samples ($S\in\mathbb{N}\text{ with }S>1$) and $N$ as the length of the RIRs ($N\in\mathbb{N}\text{ with }N>1$), and $\mathbf{y}\in\mathbb{L}^{S}\text{ where }\mathbb{L}\text{ is the set of possible labels}$. An ensemble ML model, specifically the SIREC model, is then trained using these RIRs, employing a predetermined number of estimators $N_{e}\in\mathbb{N}_{1}$, which are DTs. For each tree, two intervals are generated: a random interval, whose length varies within a adjustable maximum and minimum range across the total length of the RIRs, and a fixed interval, matching the length of the random interval but positioned at the start of the RIRs. Each random interval can be represented as a submatrix of $\mathbf{X}_{S,1:N}$ as follows in Equation (4):
$\mathbf{X}_{S,p_{i}:q_{i}}=\begin{bmatrix}v_{1,p_{i}}&v_{1,p_{i}+1}&\ldots&v_{1,q_{i}}\\ v_{2,p_{i}}&v_{2,p_{i}+1}&\ldots&v_{2,q_{i}}\\ \vdots&\vdots&\ddots&\vdots\\ v_{S,p_{i}}&v_{S,p_{i}+1}&\ldots&v_{S,q_{i}}\end{bmatrix}$ (4)
where $\mathbf{X}_{S,p_{i}:q_{i}}\in\mathbb{R}^{S\times(q_{i}-p_{i}+1)}$ with $p_{i}$ and $q_{i}$ as the start and end column indices of $\mathbf{X}_{S,1:N}$ for the $i$-th random interval ($i\in\{1,2,\ldots,N_{e}\}$), and $p_{i},q_{i}\in\{1,2,\ldots,N\}\text{ with }p_{i}<q_{i}$. For the random intervals, two feature classes are extracted. Let $\mathbf{\hat{v}}_{s_{i}}$ be defined by Equation (5) as follows:
$\mathbf{\hat{v}}_{s_{i}}=\begin{bmatrix}v_{s,p_{i}}&v_{s,p_{i}+1}&\ldots&v_{s,q_{i}}\end{bmatrix}^{T}$ (5)
where $\mathbf{\hat{v}}_{s_{i}}$ is the $s$-th row of $\mathbf{X}_{S,p_{i}:q_{i}}$. Then, $\forall s=1,2,\ldots,S:\forall i=1,2,\ldots,N_{e}$, the following transformations are applied as shown in Equations (6) and (7):
$\mathbf{\hat{x}}_{s_{i}}=\left|\text{FFT}\left(\mathbf{\hat{v}}_{s_{i}}\right)\right|$ (6)
$\hat{z}_{s_{i}}=\frac{\min(\mathbf{\hat{v}}_{s_{i}})}{\max(\mathbf{\hat{v}}_{s_{i}})}$ (7)
where $\mathbf{\hat{v}}_{s_{i}}\in\mathbb{R}^{(q_{i}-p_{i}+1)}$, $\mathbf{\hat{x}}_{s_{i}}$ are the FFT features from $\mathbf{\hat{v}}_{s_{i}}$ (the magnitude of the frequency spectrum) with $\mathbf{\hat{x}}_{s_{i}}\in\mathbb{R}^{2^{\lfloor log_{2}(q_{i}-p_{i})+0.5\rfloor}}$, and $\hat{z}_{s_{i}}$ is the min/max ratio feature from $\mathbf{\hat{v}}_{s_{i}}$ with $\hat{z}_{s_{i}}\in\mathbb{R}$. Each fixed interval can also be represented as a submatrix of $\mathbf{X}_{S,1:N}$:
$\mathbf{X}_{S,1:q_{i}-p_{i}+1}=\begin{bmatrix}v_{1,1}&v_{1,2}&\ldots&v_{1,q_{i}-p_{i}+1}\\ v_{2,1}&v_{2,2}&\ldots&v_{2,q_{i}-p_{i}+1}\\ \vdots&\vdots&\ddots&\vdots\\ v_{S,1}&v_{S,2}&\ldots&v_{S,q_{i}-p_{i}+1}\end{bmatrix}$ (8)
where $\mathbf{X}_{S,1:q_{i}-p_{i}+1}\in\mathbb{R}^{S\times(q_{i}-p_{i}+1)}$.
For the fixed intervals, two feature classes are also extracted. Let $\mathbf{v}_{s_{i}}$ be defined by Equation (9) as follows:
$\mathbf{v}_{s_{i}}=[v_{s,1},v_{s,2},\ldots,v_{s,q_{i}-p_{i}+1}]^{T}$ (9)
where $\mathbf{v}_{s_{i}}$ is the $s$-th row of $\mathbf{X}_{S,1:q_{i}-p_{i}+1}$.
Then, $\forall s=1,2,\ldots,S:\forall i=1,2,\ldots,N_{e}:$
$\mu_{s_{i}}=\frac{\sum_{k=1}^{q_{i}-p_{i}}(v_{s,k}-v_{s,k+1})}{q_{i}-p_{i}}$ (10)
$\sigma_{s_{i}}=\sqrt{\frac{1}{q_{i}-p_{i}}\sum_{k=1}^{q_{i}-p_{i}}\left(v_{s,k}-v_{s,k+1}-\mu_{s_{i}}\right)^{2}}$ (11)
where $\mathbf{v}_{s_{i}}\in\mathbb{R}^{(q_{i}-p_{i}+1)}$, $\mu_{s_{i}}$ is a specific mean feature from $\mathbf{v}_{s_{i}}$ with $\mu_{s_{i}}\in\mathbb{R}$, and $\sigma_{s_{i}}$ is a specific standard deviation feature from $\mathbf{v}_{s_{i}}$ with $\sigma_{s_{i}}\in\mathbb{R}$.
@@ II-B Training & Testing
For each DT and the associated two intervals (random & fixed), the four feature classes outlined here are extracted to train the respective tree and infer during the prediction process. The two feature classes for the fixed interval are based on novel approaches and have the following rationale: Time series data can have varying structures, and in some cases, they may exhibit symmetry along the x-axis without a vertical shift. In this project, the RIR is calculated in a manner that closely exhibits these two properties. An example of an extracted RIR is shown in Figure 4. Considering this structure of the RIRs, calculating the mean would yield a value close to zero for different classes, making it also problematic for the standard deviation, which depends on the mean. In light of this, we propose alternative calculation methods (Equation 10 & Equation 11), computing the difference between each pair of adjacent values across all values of a time series variable to determine the mean and standard deviation of the derived series, which is one value shorter than the original. This approach captures not only the magnitude of the mean and standard deviation but also the characteristic property of each RIR, combined with these two measures.
@@
Each of the $N_{\text{e}}$ trees is trained using this feature extraction methodology, forming an independent ML model that, in combination, constitutes SIREC, the complete ensemble ML model (Figure 5). In the testing process, depicted as number 3 in Figure 2, the model’s accuracy is validated using a portion of the collected RIRs and their corresponding fill level labels. Here, SIREC utilizes the intervals created during its training process, extracts relevant features, and aggregates predictions from all its DTs to derive a final result. The F1 score metric is then used to determine if the validation is sufficiently accurate to, if successful, convert the trained and validated SIREC model into C++ code. This code can then be seamlessly uploaded to the ESP8266 board using the Arduino IDE [31]–a development environment that facilitates programming and interfacing with various micro-controllers and modules. Thus, enabling local predictions using the trained ML model without the need to delegate the classification task to external resources.
Figure 5: Sonic Impulse Response Ensemble Classifier (SIREC) methodology.
III Validation
@@
III-A Performance
Data was systematically collected under varied conditions to benchmark different classification models. A selection of the characteristics of these data collection environments is presented in Table I. The first column in the dataset uniquely identifies each entry. The next five columns present the distribution of data across designated categories. The final column indicates the type of filler material used, which is either straw or cardboard (CB). A linear sine-sweep signal (Figure 3), used to derive these measurements, was played between samples 2048 and 3072 of a 4096-sample-long recording captured at a sampling rate of 10kHz. The sample interval prior to this range was utilized for noise estimation, instrumental in noise attenuation via spectral subtraction.
TABLE I: Extract of Accumulated Measurement Data
@@ III-A Performance
It is particularly noteworthy that SIREC and the Random Interval Spectral Ensemble (RISE) [34] algorithm, which secured second place in Figure 6, share two fundamental commonalities. Both methodologies are predicated on an ensemble of DTs and both engage in the generation of random intervals, within which they extract the magnitude of the frequency spectrum as one of the features. However, there are key advantages that distinguish SIREC from RISE, rendering it more beneficial. Contrary to RISE, SIREC exhibits superior speed and energy efficiency (as will be discussed in the following section and shown in Table III), as well as a significant improvement in memory efficiency. This is largely due to the fact that RISE computes the auto-correlation function, with the default setting being 100 auto-correlation terms, which inherently makes SIREC more suited for local operation on micro-controllers for classification tasks. Furthermore, SIREC tends to demonstrate marginally better average accuracy. To support this claim, Figure 7, as an extension to Figure 6, provides a detailed depiction of the F1 scores between RISE and SIREC over the 100 iterations.
@@
A significant advantage of the Random Interval Generation concept for feature extraction, as implemented in SIREC, is its inherent reduction or elimination of the need for expert knowledge about the collected data to train the model. For instance, a stochastic search across various randomly generated intervals within specified minimum and maximum lengths allows for the bypassing of such expert knowledge. This stochastic search method ensures the identification of the ideal data segment for feature extraction without necessitating extensive manual data analysis. Moreover, since SIREC is based on DTs, which represent a symbolic form of representation, acquiring expert knowledge about the data post-training is straightforward. Our stochastic search for optimal SIREC parameters (thus, the best random intervals) has achieved accuracies up to 100% in some cases. The results of this stochastic search, along with the parameters used for SIREC, are presented in Table II.
TABLE II: Stochastic Search for SIREC with Training Data $T_{1}$ $T_{2}$, Test Data $T_{3}$ $T_{4}$.
Segment Length Ne Max. Interval Length Min. Interval Length Random State F1 Score
@@
III-B Energy Measurements
In the analysis of energy consumption, our investigation reveals significant insights into the trade-offs between computational energy efficiency and predictive accuracy across three distinct algorithms: our proposed algorithm (SIREC), Time Series Forest Classifier (TSFC)[35] and RISE. The measurement method and experiment setup is based upon our previous work [36, 37] and explained in detail by [38]. The resource measurements adhere to the methodology and guidelines outlined in the Green Software Measurement Model [39]. Measurements were carried out using a computer system that included an Intel Core i5-650 CPU. The system’s memory was composed of two 2 GB RAM modules, totaling 4 GB. For storage, the computer utilized a dual-setup, comprising a 500 GB hard disk drive (HDD) and a 250 GB solid-state drive (SSD). The analysis is structured around various scenarios–configurations of segment lengths and numbers of estimators, focusing on key metrics such as energy consumption during training and testing phases, computational time, CPU usage, and F1 score. These scenarios, along with their corresponding results listed in Table III, were tested on the same datasets, $T_{1}$ and $T_{2}$ for training, and $T_{3}$ and $T_{4}$ for testing, as detailed in Table I. It is important to highlight that a random state parameter was employed across all algorithms to ensure comparability and control over experimental variability. A value of 7 was selected arbitrarily for the random state, with no intent to favor any specific algorithm, especially not SIREC.